HIGH-RESOLUTION INFRARED SPECTRA OF THE METHYL C-H-VIBRATIONS IN ETHYL RADICAL

T. HÄBER, \underline{A. C. BLAIR}, and D. J. NESBITT, JILA, University of Colorado, Boulder, Colorado 80309-0440.

All three CH-stretch vibrations of the methyl group in ethyl radical have been measured using high-resolution infrared spectroscopy in a slit supersonic discharge expansion source. Rotational assignment and analysis has been performed on each of these spectra providing rotational and centrifugal-distortion constants. We see one A-type band, which corresponds nominally to the symmetric stretch vibration, with a band origin at 2854.1027(3) cm^{−1}. Instead of a nominally degenerate vibration, the methyl asymmetric CH stretch is split into two B- and C-type bands with a separation of 124.5361 cm^{−1} due to internal rotation around the C-C bond. When compared to the splitting in nitromethane,^{a}^{b} an analogous model system for hindered internal rotation, this separation is substantially larger.

aD. C. McKean and R. A. Watt, J. Mol. Spectrosc. 61, 184 (1976)

bD. Gorse, D. Cavagnat, M. Pesquer, and C. Lapouge, J. Phys. Chem. 97, 4262 (1993)